OpenStudy (anonymous):

How do I solve an equation like this: x^2+4x+4=0 Please include steps so I can finish my homework on my own :) Thanks.

OpenStudy (anonymous):

\[x^2+4x+4=0\]

OpenStudy (anonymous):

quadratic formula :I

OpenStudy (anonymous):

you have to find discriminant at first: \[D=4^2- 4 \times 4 = 0\] When D = 0, so \[x= (-4)\div2\]x = -2

OpenStudy (anonymous):

Sorry, I'm not on discriminant yet so I have no idea what that means or how you got that :\

OpenStudy (anonymous):

x = -2, but I'm not good at explaining, sorry.

OpenStudy (anonymous):

then you can solve it by factoring (x+2)(x+2)=0 x+2=0 or x+2=0 (it's the same so the answer also is the same) x=-2

OpenStudy (anonymous):

Thanks for the help :)